OpenStudy (anonymous):

a table top is shaped like a rectangle whose perimeter is 67inches. The length is 5.5 inches more then the width. find the length and width

OpenStudy (anonymous):

w+5.5=l 2l+2w=67 2(w+5.5)+2w=67 2w+11+2w=67 4w=56 w=14 so l=19.5

OpenStudy (anonymous):

Thank you